OpenStudy (anonymous):

If 2^x=8^x-6,then what is the value of x?

hero (hero):

Hint: \(2^x = 2^{3(x-6)}\)

OpenStudy (anonymous):

8^x can be written as (2^3)^x 2x=23(x−6) as bases r equal equate powers

OpenStudy (anonymous):

Hero this Hint does not help me

hero (hero):

2nd Hint: Since the bases are the same equate the exponents, then solve for x: \(x = 3(x-6)\)

OpenStudy (anonymous):

where did you got the 3

hero (hero):

\(2^3 = 8\)

hero (hero):

Still confused?

OpenStudy (anonymous):

hold on im solving this

OpenStudy (anonymous):

8^x-6 can be written as (2^3)^x-6 (a^m )^n =a^mn

OpenStudy (anonymous):

2^x=2^3(x−6) as bases r equal equate powers x=3(x−6) n proceed

OpenStudy (anonymous):

THe answer is 9

OpenStudy (anonymous):

yes well done

OpenStudy (anonymous):

yayyy

OpenStudy (nikvist):

\[8^{x-6}\quad\mbox{or}\quad 8^{x}-6\]